OpenStudy (codyhelton):

Which of the following prime factorizations represents the greatest common factor of 162, 378, and 414? a) 2 • 3^3 b) 2 • 3^4 • 7 • 23 c) 2 • 3^2 • 7 • 23 d) 2 • 3^2

OpenStudy (codyhelton):

@Mariamikayla

OpenStudy (mariamikayla):

162|2 81|3 27|3 9|3 3|3 1 162=3^4*2 378|2 189|3 63|3 21|3 7|7 1 378=2*3^3 * 7 414|2 207|3 69|3 23|23 1 414=2* 3^2 * 23 GCF= 2* 3^2= 2* 9=18

OpenStudy (mariamikayla):

you are sure you've written corect the variants?

OpenStudy (codyhelton):

so i fixed it

OpenStudy (codyhelton):

@Mariamikayla

OpenStudy (mariamikayla):

it's variant d)